OpenStudy (anonymous):

Help with practice math questions for a competition please? (:

OpenStudy (nothingwasthesame):

ok :)

OpenStudy (anonymous):

Thanks! I have lots of different types of questions.

OpenStudy (anonymous):

Okay, here. (Sorry, stupid computer...) *If y varies directly as x and inversely as z^2, if y=6, x=3, and z=2, then which statement below is correct? A. 2xy=9z^2 B. x=2y/z^2 C. y=x+12/(z^2) D. y=8x/z^2 E. y= 72/(xz^2)